angle A and B are supplementary. Find angle A and angle B given Angle A=3x+9 and Angle B=4x-18
Answer by stanbon(75887)

angle A and B are supplementary. Find angle A and angle B given Angle A=3x+9 and Angle B=4x-18
--------------------------
If A and B are supp, A+B = 180
So, 3x+9 + 4x-18 = 180
7x -9 = 180
7x = 189
x = 27
------------
A = 3x+9 = 3*27+9 = 90
B = 4x-18 = 4*27-18 = 90
==============================
Cheers,
Stan H.